OpenStudy (anonymous):

There are 5 red and 9 green balls in a bag. If you randomly choose balls one at a time, with replacement, what is the probability of choosing 2 red balls and then 1 green ball?

OpenStudy (anonymous):

Is it .0819970845, which is 225/2744????

OpenStudy (anonymous):

\[\frac{5}{14}\times \frac{5}{14}\times \frac{9}{14}\] what ever that is

OpenStudy (anonymous):

ok, that's what i did. Thanks!!!